OpenStudy (agentjamesbond007):

Find two positive numbers whose product is 100 and whose sum is a minimum. [please explain how to solve]

myininaya (myininaya):

\[xy=100\] let's call the Sum S S=x+y

myininaya (myininaya):

now solve xy=100 for either x or y

myininaya (myininaya):

\[y=\frac{100}{x}\]

OpenStudy (anonymous):

xy = 100 or y = 100/x and we need to minimize F = x + y = x + 100/x now we are ready to take the derivative of F.

myininaya (myininaya):

\[Sum=x+\frac{100}{x}\]

myininaya (myininaya):

now do S'

OpenStudy (anonymous):

You can find the minimum of x + (100/x) using only Algebra 2 or Geometry if you're Calculus-impaired.

OpenStudy (anonymous):

Noooooo derivativeee pleaseeeeee!!!

OpenStudy (anonymous):

Simple inequalities would do just fine.

OpenStudy (anonymous):

Set S equal to the sum, x+(100/x), and solve that equation for x in terms of S. You'll get a parabola in x that has only one minimum at the Vertex of that parabola.

OpenStudy (anonymous):

If the product of the two variables is constant, the sum of the two variables will be least when they are equal.

OpenStudy (anonymous):

@FoolForMath: but how to prove that rigorously without calculus / derivatives?

OpenStudy (anonymous):

@Broken Fixer:AM-GM inequality.

OpenStudy (anonymous):

@Luis: if Sum = 100 then there is no way to attain the minimum product with positive reals (infimum is 0).